OpenStudy (anonymous):

Which of the following factors may be used to find the positive zero of the function f(x) = 2x2 + 8x - 24 ? A. (x - 4) B. (2x - 2) C. (2x - 4) D. (2x - 6)

OpenStudy (lieutenantgeneral):

2 (x - 2) (x +6) = f(x) f(x) = 2 (x^2 + 4 x - 12)